OpenStudy (anonymous):

What is the sum of the geometric sequence -4, 24, -144, ... if there are 7 terms?

OpenStudy (anonymous):

\(\LARGE S_n=\frac{a_1(1-r^n)}{1-r}\), where \(S_n\) is the sum, \(a_1\) is the term, r is the ratio and n is the number of terms. you can find the ratio by diving a term by its previous term.

OpenStudy (anonymous):

so like, -144/24?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Then once I hit the 7th term, I divide that by the previous term and I should get my answer?

OpenStudy (anonymous):

what do you mean? r= -144/24 then substitute what you get for r in the formula i typed above. you are looking for \(S_7\). you have \(a_1=-4\), \(n=7\), \(r=-144/24\)

OpenStudy (anonymous):

Oh, I get it.

OpenStudy (anonymous):

(^_^) so what did you get?

OpenStudy (anonymous):

I got -159,964

OpenStudy (anonymous):

yeah, you're right

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

no problem